OpenStudy (anonymous):

what is the answer..... (7x squared - 2x+1) + (8x cubed + 2x squared +5x -4???

OpenStudy (anonymous):

this is a two step equation

OpenStudy (drigobri):

Are you sure about this numbers? If you try to calculating (7x² - 2x + 1) = (ax² - bx +c) 1)Calculating Δ: Δ = b² - 4.a.c Δ = -2² - 4. 7. 1 Δ = 4 - 4 7. 1 Δ = -24 There are no real roots.

OpenStudy (firekat97):

I think when you open the second bracket up and collect like terms, the determinant would change and real solutions will become a possibility

OpenStudy (firekat97):

@BAdhi I can't find any zeros for this cubic :/ I tried subbing in random points as well ask factoring... any ideas?

OpenStudy (badhi):

I guess since you are asking for an answer it is an equation not jst a statement, So, \(f(x) = (7x^2-2x+1)+(8x^3+2x^2+5x-4)\) \(f(x) = 8x^3+9x^2+3x-3 \) when this is graphed a one we can see that it cuts the y = 0 axis at 0.389 so there is no way you gonna find that value by jst guessing values. for these kind of equations, there are several methods that can be followed. A famous method is the Newton-Raphson method which is an iterative method where you guess the first value which you think it is the root and with that value you can reach close to the actual root by iteration. I dont know why you say that this is two step question though : /

OpenStudy (firekat97):

@BAdhi oh okay, yeah I remember learning about that method. Thanks.